Find the power output of the engine of a 1200 kg car while the car accelerates from 30 km/h to 100 km/h in 10 s.


1
Expert's answer
2021-09-28T11:11:56-0400

The power can be found by the following equation obtained from the work-energy theorem and equation for power in terms of work and time:


"P=\frac{m}{2t}[{v_2^2-v_1^2}]=\\\space\\
=\frac{1200}{2·10}[(100·1000/3600)^2-(30·1000/3600)^2]=\\=42130\text{ W}."